On March 31, 2021, the Herzog Company purchased a factory complete with vehicles and equipment. The allocation of the total purchase price of $1,000,000 to the various types of assets along with estimated useful lives and residual values are as follows:

Asset

Cost

Estimated Residual Value

Estimated Useful Life (in years)

Land

100,000

N/A

N/A

Building

500,000

None

25

Equipment

240,000

10% of cost

8

Vehicles

160,000

12,000

8

Total

1,000,000

On June 29, 2022, equipment included in the March 31, 2021, purchase that cost $100,000 was sold for $80,000. Herzog uses the straight-line depreciation method for buildings and equipment and the double-declining-balance method for vehicles. Partial-year depreciation is calculated based on the number of months an asset is in service.

Required:

  1. Compute depreciation expense on the building, equipment, and vehicles for 202
  2. Prepare the journal entries for 2022 to record (a) depreciation on the equipment sold on June 29, 2022, and (b) the sale of the equipment. Round to the nearest whole dollar amount.
  3. Compute depreciation expense on the building, remaining equipment, and vehicles for 2022.
Answer #1

A . Depreciation expense on the building, equipment, and vehicles for 2022

Depreciation on Building

= ( (Cost -Residual value)/useful life )*9/12

=($500000-0)/25 *9/12

=Depreciation on Building = $15,000 Per Year

Depreciation on Equipment

= ( (Cost -Residual value)/useful life )*9/12

=(240000-10% of 240000)/8 *9/12

=20250 Per Year

Depreciation on Vehicle

Depreciation Rate = Depreciation rate under Straight Line * 2

= (1 / Useful life) * 2

=(1/8)*2=25%

Depreciation=Cost*Rate*9/12 months

=160000*25%*9/12

Depreciation on vehicle=$30000

B. Journal entries for 2022 to record (a) depreciation on the equipment sold on June 29, 2022

JUne 29 Depreciation expense 5625
Accumulated depreciation -equipment 5625
(Depreciation for equipment sold)
JUne 29 Cash 80000

Accumulated depreciation

loss on sale of equipment
Equipment 100000
(being sale of equipments)

Depreciation expense for 2022 = ($100,000 - $10,000) /8 * 6/12 = $5625

Depreciation expense for 2021 = ($100,000 - $10,000) /8 * 9/12 = $ 8438

Total Accumulated depreciation = $ 8437.5 + $ 5625 = $ 14063

Loss on sale = $100,000 - $80,000 - $ 14062.5 = $ 59378

C. Depreciation expense on the building, remaining equipment, and vehicles for 2022.

Depreciation on Building =
Depreciation for year 2021 = ($500,000 - 0) /25 = $20000

Depreciation on Equipment =
Residual value = ($240,000 - $100,000) * 10% = $14000
Depreciation for year 2022 - ($140000 - $14000) /8 = $ 15750

Depreciation on Vehicles :
Book Value at end of 2021 = $160,000 - $30,000 = $130,000
Depreciation for year 2022 = $130,000 * 25% = $32500